# Quantum backreaction in an analogue black hole

### v1

@@ IV.1.1 The three dimensional case
In 3 dimensions, the last term of the right hand side of expression (62) has an utraviolet divergence, associated to the anomalous average $\langle\hat{a}\hat{a}\rangle$ in (45). But, expanding the coupling constant to second order in the scattering length regularizes this divergence (see e.g., [45, 46]), leading to the Lee-Huang-Yang expression [47, 7]
@@ IV.1.2 The one dimensional case
We are mainly interested in 1D configurations where the situation is actually simpler because in this case $g^{(2)}$ given from the Bogoliubov expression (45) is regular. We get [48]
@@
In this expression we replaced in the corrective term the Gross-Pitaevskii healing length $\xi_{\rm\scriptscriptstyle GP}$ by $\xi$, which is legitimate. The result (66) has already been obtained in Ref. [30] and corresponds to the weak interaction expansion of the exact Lieb-Liniger result [49]. Inserted in the thermodynamic relation $mc^{2}=\rho(\partial\mu/\partial\rho)$ it yields
@@
This formula defines the speed of sound $c$ as a function of the density $\rho$ in a 1D homogeneous system at zero temperature. It corrects the Gross-Pitaevskii formula $c_{\rm\scriptscriptstyle GP}=(g\rho_{\rm\scriptscriptstyle GP}/m)^{1/2}$. It is clear from (66) that in 1D the dimensionless small parameter of our approach is $(\rho\xi)^{-1}$, implying that our results are expected to be valid only in the limit $\rho\xi\gg 1$. However, it was observed by Lieb [50] that expression (67) agrees very well with the exact result down to $\rho\xi\sim 0.3$. In the case of interest for us, a typical experimental order of magnitude is $\rho\xi\approx 30\leftrightarrow 60$ [51], i.e., quite far from the regime where (67) becomes incorrect.
@@ IV.2 Analogous black hole in a 1D stationary transonic flow
The asymptotic expressions (73) and (76) show that, whereas the upstream density and velocity modifications tend to constant values, the downstream asymptotic profile typically displays undulations with a wave vector $\kappa_{d}$ associated to a zero energy channel (identified in Fig. 3 in Appendix B). Such stationary undulations have been predicted to appear in the supersonic region of a white hole configuration non-linearly stimulated by an external seed impinging from the subsonic region [52]. A similar stimulated process is not possible in a black hole configuration because the group velocity of the corresponding zero mode is directed outward the supersonic region [^2] In the terminology of Appendix B, the zero energy channel is not outgoing, see Fig. 3., and thus cannot be excited by a source located outside the black hole. In our case, although we consider a black hole configuration, the appearance of undulations is however perfectly legitimate because it results from a spontaneous process, for which the effective “seed” (the source term $S_{d}$ which accounts for quantum fluctuations) exists throughout the whole supersonic region.
The precise asymptotic behavior of $\delta\rho(x)$ and $\delta V(x)$ depends on the constants ${\cal A}_{\alpha}$ and $\phi$ in (73) and (76) which can only be determined by a full solution of the back reaction equations (68), (69). This, in turn, necessitates a determination of the source terms in the whole physical space, which can be accurately done only by correctly taking account of zero-mode solutions of the Bogoliubov equations [30]. It has been shown in Ref. [53] that this is indeed crucial for accurately determining the local density-density correlation function $g^{(2)}(x)$ in the vicinity of the horizon. However, the contributions of the zero-modes are not necessary when considering the upstream and downstream asymptotic regions, far away from the horizon. This noticeably simplifies the computation of the contribution of the quantum fluctuations to the asymptotic local density-density correlation function $g_{\alpha}^{(2)}$ and to the asymptotic current $j_{\alpha}$. We show in Appendix C that they are of the form
@@
In the present pilot study we focus on the mean values $\delta\rho_{\alpha}$ and $\delta V_{\alpha}$ of the asymptotic behaviors (73) and (76) and defer a full numerical solution of the back reaction equations to a future work. It should be kept in mind that, because of the downstream undulations, the modification $\delta\rho_{d}$ and $\delta V_{d}$ are average quantities. The situation is simpler in the upstream region where the asymptotic profile is flat. Hence, it is more appropriate to solve Eqs. (74) and (77) fixing boundary conditions in the upstream region where the modified flow pattern is asymptotically featureless (constant density and velocity). In an equilibrium situation, in the absence of background flow velocity, it was argued in Ref. [30] that the appropriate condition is $\delta\mu=0$, which corresponds to a grand canonical situation. In an homogeneous system, in the presence of a constant background flow, Galilean invariance furthermore imposes that the beyond mean-field effects do not modify the flow velocity. [^3] Consider a uniform flow of constant density and velocity in the laboratory frame. In the comoving frame of the fluid, the Gross–Pitaevskii background velocity vanishes. Since backreaction —i.e., beyond mean-field corrections— cannot generate a spontaneous flow in this frame, it follows that in the laboratory frame the initial velocity remains unaffected by backreaction. Because of the asymptotic homogeneity of the upstream flow we thus impose the similar condition $\delta V_{u}=0$. From Eqs. (72), (74) and (77) these two conditions determine $\delta J$, $\delta\rho_{u}$, $\delta\rho_{d}$ and $\delta V_{d}$.
@@
In the framework of the Gross-Pitaevskii approximation, it has been shown that the main characteristics of the analog black hole and of the associated Hawking radiation are determined by the values of the asymptotic Mach numbers $M_{\scriptscriptstyle\rm GP}^{\alpha}=V_{\scriptscriptstyle\rm GP}^{\alpha}/c_{\scriptscriptstyle\rm GP}^{\alpha}$ (see, e.g., Ref. [54]). It is thus important to evaluate to which extent these quantities are affected by the modifications $\delta\rho_{\alpha}$ and $\delta V_{\alpha}$ of the asymptotic densities. backreaction effects modify the Gross-Pitaevskii result to $M_{\alpha}=M^{\alpha}_{\rm\scriptscriptstyle GP}+\delta M_{\alpha}$ with
@@
where use has been made of expression (67) for the one-dimensional speed of sound. In the upper panel of Figs. 1 we display the relative modifications $\delta M_{\alpha}/M_{\scriptscriptstyle\rm GP}^{\alpha}$ for the waterfall configuration (the results for other configurations are presented in Appendix E). In region $\alpha$ ($\alpha=u$ or $d$) this modification is proportional to the small parameter $(\rho_{\scriptscriptstyle\rm GP}^{\alpha}\xi_{\scriptscriptstyle\rm GP}^{\alpha})^{-1}$, the value of which depends of the experimental situation, and is typically of order 0.02 in Steinhauer’s experiment [51]. This is the reason why $\delta M_{\alpha}/M_{\scriptscriptstyle\rm GP}^{\alpha}$ is rescaled by $\rho_{\scriptscriptstyle\rm GP}^{\alpha}\xi_{\scriptscriptstyle\rm GP}^{\alpha}$ in the upper panel of this figure. In the same plot, the result obtained when discarding the contribution of Hawking radiation [i.e., by removing the contributions ${\cal G}_{\alpha}^{(H)}$ and ${\cal J}_{\alpha}^{(H)}$ in Eqs. (78)] are represented by dashed lines. The blue horizontal dashed line corresponds to the value $1/4\pi$ which would be the rescaled relative modification of the Gross-Pitaevskii Mach number of a homogeneous condensate (constant density $\rho_{\scriptscriptstyle\rm GP}^{u}$ and constant velocity $V_{\scriptscriptstyle\rm GP}^{u}$) induced by beyond mean-field effects. Concerning the downstream modification $\delta M_{d}$ of the Mach number, we stress that it is evaluated from the average quantities $\delta V_{d}$ and $\delta\rho_{d}$. When the period $2\pi/\kappa_{d}$ of the undulations of the downstream density is small compared to $\xi_{\scriptscriptstyle\rm GP}^{d}$ (i.e., when $M_{\scriptscriptstyle\rm GP}^{d}\gg 1$) it is a priori not possible to define a local speed of sound: in this limit $\delta M_{d}$ defined by (79) is not the spatial average of an hypothetical local downstream Mach number.
@@
The primary observation drawn from the upper panel of Figure 1 is that the relative modifications in $M_{u}$ and $M_{d}$ are small. The upstream Mach number weakly decreases and the downstream one weakly increases. It thus appears that, for the chosen boundary conditions, the transonic character of the flow is not modified by quantum fluctuations and backreaction. The flows are poorly affected in the whole range $0<M_{\scriptscriptstyle\rm GP}^{u}<1$, which legitimates our perturbative approach. Another clear feature is that the solid lines depart weakly from the dashed ones. This means that Hawking radiation has a lesser impact on the Mach number than the standard beyond mean-field quantum fluctuations. However, one should keep in mind that the beyond mean-field corrections are present even in the absence of an acoustic horizon, i.e., they should be taken into account already in an homogeneous system, before the formation of the analog black hole. It is thus appropriate to distinguish the beyond-mean field terms from the backreaction effects truly induced by Hawking radiation. This is the reason why we represent in the lower panel of Fig. 1 the quantities $\delta M_{\alpha}^{(H)}$ ($\alpha=u$ or $d$) which are the difference between the modification of the Mach number computed with and without the Hawking contribution. For being able to compare the relative upstream and downstream influence of Hawking radiation, in the lower panels we rescale $\delta M_{u}^{(H)}/M_{\scriptscriptstyle\rm GP}^{u}$ and $\delta M_{d}^{(H)}/M_{\scriptscriptstyle\rm GP}^{d}$ by the same quantity $\rho_{\scriptscriptstyle\rm GP}^{u}\xi_{\scriptscriptstyle\rm GP}^{u}$. It appears that Hawking radiation induces positive modification of the upstream and downstream Mach numbers, of roughly the same order. We show in Appendix E that, while the positivity of $\delta M_{u}^{(H)}$ observed in the lower panel of Fig. 1 is a general feature due to our specific choice of boundary conditions, $\delta M_{d}^{(H)}$ may be positive or negative in other configurations.
@@ V Conclusion and perspectives
Our main interest lies in 1D situations with a transonic flow realizing a sonic horizon. In this case, we solved the stationary backreaction equations far from the acoustic horizon (deep in the upstream and downstream regions). Whereas the asymptotic upstream flow tends to a homogeneous limit, the existence of downstream channels of zero energy (themselves resulting from the transonic character of the flow) induces, through backreaction effects, stationary density and velocity undulations in the downstream region. The existence of undulations in the interior of the analogue black hole is a nonlinear effect of quantum backreaction, not expected within a Gross-Pitaevskii approach.
A natural extension of our work is to consider a generic situation, such as the ones experimentally realized in Refs. [51, 3], and to determine a stationary solution of the backreaction equations, not only in the asymptotic regions, but in the whole physical space. In this case, the theoretical expansion of the quantum fluctuations should explicitly account for the existence of zero-modes of Bogoliubov equations [55]. In such generic situations, on the basis of the orders of magnitudes we obtained in the asymptotic regions, we expect that the backreaction effects should be small, and would not drastically affect the leading order Gross-Pitaevskii flow profile.
It should be noted that the detailed form of the stationary solutions and of the modifications of the upstream and downstream Mach numbers depend on the boundary conditions imposed at infinity. Simple physical arguments lead to impose $\delta\mu=0=\delta V_{u}$, but other more elaborate conditions may be imposed which modify the specifics of the results presented in Figs. 1, 4, 5 and 6 (such as the positivity of $\delta M_{u}^{(H)}$ for instance). Since the backreaction equations we derived are valid for a time-dependent background flow, a natural way to circumvent this issue would be to study the dynamics of formation of an acoustic horizon in a quasi 1D BEC, as already studied in Refs. [56, 57, 58, 59, 60, 61, 62], with account of quantum backreaction effects. Work in this direction is in progress.
@@ Appendix A Different black-hole configurations
In this study, we examine three different types of analogue black holes, referred to as “waterall”, $\delta$-peak” and “flat profile” in Ref. [54]. They all correspond to the flow of a one-dimensional BEC which is subsonic in the upstream region and supersonic downstream. The associated order parameter is a stationary solution of the Gross-Pitaevskii equation (18). According to the convention adopted in this work, the corresponding fields and characteristic quantities should be denoted with a subscript “GP”. However, to facilitate readability, we omit this index in this appendix.
@@
The analog black hole realized in the 2019 Technion experiment [3] is close to the waterfall configuration with $M_{d}=2.9$, cf. the discussion in [53].
@@
The flat profile configuration, introduced in Refs. [4, 56] is a idealized setting in which the Gross-Pitaevskii density and velocity are constant: $\rho(x)=\rho_{0}$ and $V(x)=V_{0}$. An acoustic horizon can be implemented in such a configuration by means of a step-like nonlinear constant $g(x)$ combined with a step-like external potential $U(x)$:

### v2

@@ IV.1.1 The three dimensional case
In 3 dimensions, the last term of the right hand side of expression (63) has an utraviolet divergence, associated to the anomalous average $\langle\hat{a}\hat{a}\rangle$ in (46). But, expanding the coupling constant to second order in the scattering length regularizes this divergence (see e.g., [Pitaevskii2016, Pethick2011]), leading to the Lee-Huang-Yang expression [Lee1957, Beliaev1958b]
@@ IV.1.2 The one dimensional case
We are mainly interested in 1D configurations where the situation is actually simpler because in this case $g^{(2)}$ given from the Bogoliubov expression (46) is regular. We get [Gangardt2003]
@@
In this expression we replaced in the corrective term the Gross-Pitaevskii healing length $\xi_{\rm\scriptscriptstyle GP}$ by $\xi$, which is legitimate. The result (67) has already been obtained in Ref. [Mora2003] and corresponds to the weak interaction expansion of the exact Lieb-Liniger result [Lieb1963a]. Inserted in the thermodynamic relation $mc^{2}=\rho(\partial\mu/\partial\rho)$ it yields
@@
This formula defines the speed of sound $c$ as a function of the density $\rho$ in a 1D homogeneous system at zero temperature. It corrects the Gross-Pitaevskii formula $c_{\rm\scriptscriptstyle GP}=(g\rho_{\rm\scriptscriptstyle GP}/m)^{1/2}$. It is clear from (67) that in 1D the dimensionless small parameter of our approach is $(\rho\xi)^{-1}$, implying that our results are expected to be valid only in the limit $\rho\xi\gg 1$. However, it was observed by Lieb [Lieb1963b] that expression (68) agrees very well with the exact result down to $\rho\xi\sim 0.3$. In the case of interest for us, a typical experimental order of magnitude is $\rho\xi\approx 30\leftrightarrow 60$ [Steinhauer2016], i.e., quite far from the regime where (68) becomes incorrect.
@@ IV.2 Analogous black hole in a 1D stationary transonic flow
The asymptotic expressions (74) and (77) show that, whereas the upstream density and velocity modifications tend to constant values, the downstream asymptotic profile typically displays undulations with a wave vector $\kappa_{d}$ associated to a zero energy channel (identified in Fig. 3 in Appendix B). Such stationary undulations have been predicted to appear in the supersonic region of a white hole configuration non-linearly stimulated by an external seed impinging from the subsonic region [Mayoral2011]. A similar stimulated process is not possible in a black hole configuration because the group velocity of the corresponding zero mode is directed outward the supersonic region [^2] In the terminology of Appendix B, the zero energy channel is not outgoing, see Fig. 3., and thus cannot be excited by a source located outside the black hole. In our case, although we consider a black hole configuration, the appearance of undulations is however perfectly legitimate because it results from a spontaneous process, for which the effective “seed” (the source term $S_{d}$ which accounts for quantum fluctuations) exists throughout the whole supersonic region.
The precise asymptotic behavior of $\delta\rho(x)$ and $\delta V(x)$ depends on the constants ${\cal A}_{\alpha}$ and $\phi$ in (74) and (77) which can only be determined by a full solution of the backreaction equations (69), (70). This, in turn, necessitates a determination of the source terms in the whole physical space, which can be accurately done only by correctly taking account of zero-mode solutions of the Bogoliubov equations [Mora2003]. It has been shown in Ref. [Isoard2020] that this is indeed crucial for accurately determining the local density-density correlation function $g^{(2)}(x)$ in the vicinity of the horizon. However, the contributions of the zero-modes are not necessary when considering the upstream and downstream asymptotic regions, far away from the horizon. This noticeably simplifies the computation of the contribution of the quantum fluctuations to the asymptotic local density-density correlation function $g_{\alpha}^{(2)}$ and to the asymptotic current $j_{\alpha}$. We show in Appendix C that they are of the form
@@
In the present pilot study we focus on the mean values $\delta\rho_{\alpha}$ and $\delta V_{\alpha}$ of the asymptotic behaviors (74) and (77) and defer a full numerical solution of the backreaction equations to a future work. It should be kept in mind that, because of the downstream undulations, the modification $\delta\rho_{d}$ and $\delta V_{d}$ are average quantities. The situation is simpler in the upstream region where the asymptotic profile is flat. Hence, it is more appropriate to solve Eqs. (75) and (78) fixing boundary conditions in the upstream region where the modified flow pattern is asymptotically featureless (constant density and velocity). In an equilibrium situation, in the absence of background flow velocity, it was argued in Ref. [Mora2003] that the appropriate condition is $\delta\mu=0$, which corresponds to a grand canonical situation. In an homogeneous system, in the presence of a constant background flow, Galilean invariance furthermore imposes that the beyond mean-field effects do not modify the flow velocity. [^3] Consider a uniform flow of constant density and velocity in the laboratory frame. In the comoving frame of the fluid, the Gross–Pitaevskii background velocity vanishes. Since backreaction —i.e., beyond mean-field corrections— cannot generate a spontaneous flow in this frame, it follows that in the laboratory frame the initial velocity remains unaffected by backreaction. Because of the asymptotic homogeneity of the upstream flow we thus impose the similar condition $\delta V_{u}=0$. From Eqs. (73), (75) and (78) these two conditions determine $\delta J$, $\delta\rho_{u}$, $\delta\rho_{d}$ and $\delta V_{d}$.
@@
In the framework of the Gross-Pitaevskii approximation, it has been shown that the main characteristics of the analog black hole and of the associated Hawking radiation are determined by the values of the asymptotic Mach numbers $M_{\scriptscriptstyle\rm GP}^{\alpha}=V_{\scriptscriptstyle\rm GP}^{\alpha}/c_{\scriptscriptstyle\rm GP}^{\alpha}$ (see, e.g., Ref. [Larre2012]). It is thus important to evaluate to which extent these quantities are affected by the modifications $\delta\rho_{\alpha}$ and $\delta V_{\alpha}$ of the asymptotic densities. Backreaction effects modify the Gross-Pitaevskii result to $M_{\alpha}=M^{\alpha}_{\rm\scriptscriptstyle GP}+\delta M_{\alpha}$ with
@@
where use has been made of expression (68) for the one-dimensional speed of sound. In the upper panel of Figs. 1 we display the relative modifications $\delta M_{\alpha}/M_{\scriptscriptstyle\rm GP}^{\alpha}$ for the waterfall configuration (the results for other configurations are presented in Appendix E). In region $\alpha$ ($\alpha=u$ or $d$) this modification is proportional to the small parameter $(\rho_{\scriptscriptstyle\rm GP}^{\alpha}\xi_{\scriptscriptstyle\rm GP}^{\alpha})^{-1}$, the value of which depends of the experimental situation, and is typically of order 0.02 in Steinhauer’s experiment [Steinhauer2016]. This is the reason why $\delta M_{\alpha}/M_{\scriptscriptstyle\rm GP}^{\alpha}$ is rescaled by $\rho_{\scriptscriptstyle\rm GP}^{\alpha}\xi_{\scriptscriptstyle\rm GP}^{\alpha}$ in the upper panel of this figure. In the same plot, the result obtained when discarding the contribution of Hawking radiation [i.e., by removing the contributions ${\cal G}_{\alpha}^{(H)}$ and ${\cal J}_{\alpha}^{(H)}$ in Eqs. (79)] are represented by dashed lines. The blue horizontal dashed line corresponds to the value $1/4\pi$ which would be the rescaled relative modification of the Gross-Pitaevskii Mach number of a homogeneous condensate (constant density $\rho_{\scriptscriptstyle\rm GP}^{u}$ and constant velocity $V_{\scriptscriptstyle\rm GP}^{u}$) induced by beyond mean-field effects. Concerning the downstream modification $\delta M_{d}$ of the Mach number, we stress that it is evaluated from the average quantities $\delta V_{d}$ and $\delta\rho_{d}$. When the period $2\pi/\kappa_{d}$ of the undulations of the downstream density is small compared to $\xi_{\scriptscriptstyle\rm GP}^{d}$ (i.e., when $M_{\scriptscriptstyle\rm GP}^{d}\gg 1$) it is a priori not possible to define a local speed of sound: in this limit $\delta M_{d}$ defined by (80) is not the spatial average of an hypothetical local downstream Mach number.
@@
The primary observation drawn from the upper panel of Figure 1 is that the relative modifications in $M_{u}$ and $M_{d}$ are small. The upstream Mach number weakly decreases and the downstream one weakly increases. It thus appears that, for the chosen boundary conditions, the transonic character of the flow is not modified by quantum fluctuations and backreaction. The flows are poorly affected in the whole range $0<M_{\scriptscriptstyle\rm GP}^{u}<1$, which legitimates our perturbative approach. Another clear feature is that the solid lines depart weakly from the dashed ones. This means that Hawking radiation has a lesser impact on the Mach number than the standard beyond mean-field quantum fluctuations. However, one should keep in mind that the beyond mean-field corrections are present even in the absence of an acoustic horizon, i.e., they should be taken into account already in a homogeneous system, before the formation of the analog black hole. It is thus appropriate to distinguish the beyond-mean field terms from the backreaction effects truly induced by Hawking radiation. This is the reason why we represent in the lower panel of Fig. 1 the quantities $\delta M_{\alpha}^{(H)}$ ($\alpha=u$ or $d$) which are the difference between the modification of the Mach number computed with and without the Hawking contribution. For being able to compare the relative upstream and downstream influence of Hawking radiation, in the lower panels we rescale $\delta M_{u}^{(H)}/M_{\scriptscriptstyle\rm GP}^{u}$ and $\delta M_{d}^{(H)}/M_{\scriptscriptstyle\rm GP}^{d}$ by the same quantity $\rho_{\scriptscriptstyle\rm GP}^{u}\xi_{\scriptscriptstyle\rm GP}^{u}$. It appears that Hawking radiation induces positive modification of the upstream and downstream Mach numbers, of roughly the same order. We show in Appendix E that, while the positivity of $\delta M_{u}^{(H)}$ observed in the lower panel of Fig. 1 is a general feature due to our specific choice of boundary conditions, $\delta M_{d}^{(H)}$ may be positive or negative in other configurations.
@@ V Conclusion and perspectives
Our main interest lies in 1D situations with a transonic flow realizing a sonic horizon. In this case, we solved the stationary backreaction equations far from the acoustic horizon (deep in the upstream and downstream regions). Whereas the asymptotic upstream flow tends to a homogeneous limit, the existence of downstream channels of zero energy (themselves resulting from the transonic character of the flow) induces, through backreaction effects, stationary density and velocity undulations in the downstream region. The existence of undulations in the interior of the analog black hole is a nonlinear effect of quantum backreaction, not expected within a Gross-Pitaevskii approach.
A natural extension of our work is to consider a generic situation, such as the ones experimentally realized in Refs. [Steinhauer2016, deNova2019], and to determine a stationary solution of the backreaction equations, not only in the asymptotic regions, but in the whole physical space. In this case, the theoretical expansion of the quantum fluctuations should explicitly account for the existence of zero-modes of Bogoliubov equations [Blaizot1986]. In such generic situations, on the basis of the orders of magnitudes we obtained in the asymptotic regions, we expect that the backreaction effects should be small, and would not drastically affect the leading order Gross-Pitaevskii flow profile.
It should be noted that the detailed form of the stationary solutions and of the modifications of the upstream and downstream Mach numbers depend on the boundary conditions imposed at infinity. Simple physical arguments lead to impose $\delta\mu=0=\delta V_{u}$, but other more elaborate conditions may be imposed which modify the specifics of the results presented in Figs. 1, 4, 5 and 6 (such as the positivity of $\delta M_{u}^{(H)}$ for instance). Since the backreaction equations we derived are valid for a time-dependent background flow, a natural way to circumvent this issue would be to study the dynamics of formation of an acoustic horizon in a quasi 1D BEC, as already studied in Refs. [Carusotto2008, Kamchatnov2012, Michel2015, Wang2017, Tettamanti2020, Kolobov2021, Fabbri2021], with account of quantum backreaction effects. Work in this direction is in progress.
@@ Appendix A Different black-hole configurations
In this study, we examine three different types of analog black holes, referred to as “waterall”, $\delta$-peak” and “flat profile” in Ref. [Larre2012]. They all correspond to the flow of a one-dimensional BEC which is subsonic in the upstream region and supersonic downstream. The associated order parameter is a stationary solution of the Gross-Pitaevskii equation (19). According to the convention adopted in this work, the corresponding fields and characteristic quantities should be denoted with a subscript “GP”. However, to facilitate readability, we omit this index in this appendix.
@@
The analog black hole realized in the 2019 Technion experiment [deNova2019] is close to the waterfall configuration with $M_{d}=2.9$, cf. the discussion in [Isoard2020].
@@
The flat profile configuration, introduced in Refs. [Balbinot2008, Carusotto2008] is a idealized setting in which the Gross-Pitaevskii density and velocity are constant: $\rho(x)=\rho_{0}$ and $V(x)=V_{0}$. An acoustic horizon can be implemented in such a configuration by means of a step-like nonlinear constant $g(x)$ combined with a step-like external potential $U(x)$: